# Engineering RAG Systems for Real-World Applications: Design, Development, and Evaluation

### v1

@@ Abstract
Index Terms:
Empirical Software Engineering, AI System Lifecycle, Generative AI, RAG, LLMs, System Design, System Implementation, Human Centered Evaluation
This is the author’s preprint version of a paper accepted to the 51st Euromicro Conference Series on Software Engineering and Advanced Applications (SEAA 2025). The final version will be published by IEEE.
I Introduction
Retrieval-Augmented Generation (RAG) has the capability to enhance Large Language Models (LLMs) by retrieving relevant external knowledge, thereby improving the accuracy of Generative AI (GenAI) applications. While GenAI has seen use in software engineering [1], RAG extends its value to broader domains by combining parametric and non-parametric memory, effectively addressing the limitations of static knowledge bases [2]. Recent advances in retrieval-augmented LLMs enable real-time information retrieval, reducing hallucinations and improving response reliability [3].
The foundational work by Lewis et al. [4] established RAG as a standard for tasks like question answering and knowledge retrieval. Early frameworks such as REALM and RAG demonstrated the benefits of combining dense retrieval with language generation for open-domain tasks [5]. However, most later research has focused on improving retrieval architectures and reducing hallucinations, often evaluated only on clean, English-centric benchmarks [6]. There remains limited exploration of domain-specific, multilingual, or real-world deployments, which this paper addresses through the development and evaluation of RAG systems across diverse application settings, with a focus on retrieval quality and system design.
@@ I Introduction
This study investigates how RAG systems can be engineered and evaluated in real-world contexts. It emphasizes system design and development for domain-specific applications, user evaluation across criteria such as ease of use and relevance, and lessons that inform future engineering practices. Guided by these objectives, this paper addresses the following three Research Questions (RQs):
  • RQ1: How can RAG systems be designed and developed to address real-world system needs across diverse application domains?
@@
  • RQ3: What are the lessons learned from engineering RAG systems for real-world applications?
To address these research questions, we developed five domain-specific RAG systems and evaluated them through a web-based user study with 100 participants. The evaluation focused on usability, retrieval relevance, transparency, and other user-centered factors. Full methodological details are provided in Section III.
The contributions of this paper are as follows:
  • End-to-end development and deployment of RAG systems for multilingual, domain-specific applications.
  • User-centered evaluation demonstrating real-world performance across usability and accuracy metrics.
  • Practical engineering insights to guide the design of reliable and maintainable RAG pipelines.
  • System-level considerations for integrating RAG into real-world AI-based software, contributing to software engineering practice.
Paper Structure: Section II reviews related work on RAG and its applications. Section III describes the study design. Section IV explains the implementation of five real-world RAG systems. Section V presents user system evaluation, and Section VI outlines key lessons learned. Section VII discusses study limitations, and Section VIII concludes with future directions.
II Related Work
Retrieval-Augmented Generation (RAG) improves the factual accuracy and contextual relevance of Large Language Models (LLMs) by incorporating real-time external information, making it especially valuable for complex tasks such as question answering, legal reasoning, and summarization [7]. Recent work has demonstrated RAG’s utility in taxonomy-driven dataset design [8], token-efficient document handling [7], and multimodal applications that combine text and images via Vision-Language Models (VLMs) such as VISRAG [9]. Despite these advances, OCR noise remains a limiting factor in retrieval fidelity [10]. Ongoing research addresses this by refining dataset construction [8], tackling architectural scalability [11], improving query-document alignment through prompt engineering [12], and applying speculative retrieval to boost performance in multimodal settings [13].
RAG has been widely applied in software engineering to support code understanding and developer tasks. StackRAG [14] leverages Stack Overflow content to enhance developer assistance, while CodeQA [15] employs LLM agents with retrieval augmentation for programming queries. Ask-EDA [16] addresses hallucination reduction in Electronic Design Automation (EDA) via hybrid retrieval. In industrial contexts, Khan et al. [17] examine PDF-focused retrieval challenges, and Xiaohua et al. [18] propose re-ranking and repacking strategies for pipeline optimization. In healthcare, MEDGPT [19] extracts structured insights from PubMed to support diagnostics, while Path-RAG [20] improves pathology image retrieval with knowledge-guided methods. Alam et al. [21] introduce a multi-agent retriever for radiology reports, enhancing clinical transparency, and Guo et al. [22] present LightRAG, a graph-based retriever that boosts knowledge precision across medical domains.
@@ II Related Work
The rise of autonomous AI agents has further improved Retrieval-Augmented Generation (RAG) by enabling self-directed reasoning, adaptive retrieval, and memory persistence. Wang et al. [28] survey LLM-driven agent architectures, while Liu et al. [29] benchmark multi-turn reasoning through AgentBench. AgentTuning by Zeng et al. [30] enhances instruction tuning for retrieval-based decisions. Singh et al. [31] categorize Agentic RAG into single-agent, multi-agent, and graph-based designs, highlighting dynamic tool use. On the retrieval side, Yan et al. [32] introduce CRAG to reduce hallucinations using confidence-based filtering, and Li et al. [33] improve precision through contrastive in-context learning and focus-mode filtering—strengthening RAG’s reliability in complex scenarios.
Conclusive Summary: While RAG continues to advance, challenges in retrieval accuracy, response reliability, and scalability remain [34]. Although hybrid strategies [24], autonomous agents [31], and correction techniques [32, 33, 26] have been explored, there is still limited evaluation in domain-specific settings. This paper addresses that gap by implementing and assessing five RAG systems across key sectors, offering practical insights into their real-world performance and deployment potential.
III Study Design
Figure 1 provides an overview of the methodological steps, from pipeline development to user-centered evaluation.
@@ III Study Design
To investigate the design and real-world performance of RAG-based systems, we implemented five optimized pipelines across distinct domains and conducted a structured user evaluation to assess their effectiveness and user reception.
III-A Implementing RAG Systems
@@ III-A1 Domain Selection
III-A2 System Design
Figure 2: Design overview of the RAG systems developed for five domain-specific applications.
@@ III-A3 Core Components
III-B System Evaluation Method
@@
After using the systems, participants completed a standardized survey covering six criteria: Ease of Use, Relevance of Info, Transparency, System Responsiveness, Accuracy of Answers, and Recommendation. The survey included both Likert-scale questions and open-ended feedback. This approach provided both quantitative ratings and qualitative insights into system performance. We reviewed the open-ended feedback to identify common themes in participants’ experiences. We also referred to development notes taken throughout the project. These helped us recognize recurring issues and informed the lessons described in Section VI.
IV Systems Implementation
This section outlines the end-to-end implementation of five RAG-based systems designed for real-world deployment across diverse domains. Each system was developed to address domain-specific retrieval challenges by integrating embedding models, vector databases, and LLMs. The implementations varied based on task complexity, document type, language requirements, and deployment constraints, demonstrating the adaptability of RAG pipelines in applied settings.
@@ IV Systems Implementation
    Kankaanpää City AI: This system enhances transparency of government records. It processes over 1,000 PDFs from 2023–2024, indexing them in FAISS for accurate retrieval of policy documents. The system uses text-embedding-ada-002 as the embedding model to convert documents into vector representations, and gpt-4o-mini as the LLM to generate context-aware responses. This setup allows users to search and access municipal decisions, infrastructure projects, and public policies with ease.
@@
    Disarm RAG: It is designed to deliver real-time insights into cyber threats, and forensic investigations. It is hosted on a secure server at CSC [^6] https://research.csc.fi/cloud-computing/ (Finnish IT Center for Science), ensuring full data privacy, and uses LLaMA 2-uncensored via Ollama to enable open access to cybersecurity knowledge. The system integrates red team techniques (e.g., phishing, deepfake disinformation, privilege escalation) and blue team strategies (e.g., bot detection, misinformation control, network forensics), grounded in the Disarm Framework. It supports queries such as “How would you create a deepfake to discredit a public figure?” and “What are the latest techniques for bypassing multi-factor authentication (MFA)?”, as well as defensive questions like “How would you detect a disinformation campaign early on?” and “What are effective countermeasures against deepfake-based phishing attacks?”.
@@
All systems developed in this study comply with GDPR standards to ensure responsible handling of user interactions and system outputs. To support transparency, most systems display source references alongside AI-generated responses. An exception is Disarm RAG, where source citations are omitted due to cybersecurity sensitivity.
V Systems Evaluation
Understanding the real-world effectiveness of RAG-based systems requires moving beyond technical benchmarks to incorporate user-centered evaluation. We conducted a structured user study across five domain-specific deployments, capturing both system performance metrics and user perceptions of trust, relevance, and usability. This practical feedback offers a grounded view of system behavior in real settings and highlights opportunities for targeted improvements.
V-A Participant Demographics and RAG Orientation
@@
Figure 3: Participant profiles and their interaction with RAG systems.
@@
    AI-Generated vs. Manual Document Search: Participants exhibited a task-sensitive perspective on AI assistance. While a substantial majority (83%) preferred AI-generated responses depending on the nature of the task, only (9%) expressed a consistent preference for AI over manual methods. Conversely, (8%) favored manual search regardless of context. These findings suggest that trust in RAG systems is not absolute but contingent—underscoring the importance of response relevance, transparency, and alignment with user intent.
@@
V-B Survey Instrument and Case-wise Findings
@@
VI Lessons Learned
While developing and evaluating the five RAG systems, we encountered several technical, operational, and ethical challenges. Based on what we observed during implementation and what participants shared in the evaluation, we summarized a set of lessons that reflect the most common and recurring issues across domains.
VI-A Technical Development
@@
  • FAISS Scalability Hits Limits: With large corpora (>>>10k embeddings), FAISS latency increased noticeably. Metadata filtering by document type reduced search time.
@@ VI-B Operational Factors
  • SQLite for Tracking User Interaction: We used SQLite to log user questions, responses, and ratings (e.g., in AgriHubi). This lightweight store helped identify system failures and understand user behavior.
@@ VI-C Ethical Considerations
VII Study Limitations
This study presents findings grounded in the design, deployment, and evaluation of five domain-specific RAG systems, but several limitations must be acknowledged. First, while our evaluation involved 100 participants across diverse roles including researchers, practitioners, and domain experts, approximately 20% of the sample consisted of students. Although these students had relevant technical or domain experience, their feedback may reflect differing expectations or usage behaviour compared to full-time professionals. This demographic distribution, while broad, could influence the generalizability of findings to strictly industrial settings.

### v2

@@ Abstract
1 Introduction
Retrieval-Augmented Generation (RAG) is an approach that integrates external knowledge retrieval into language model outputs to improve accuracy and relevance. RAG enhances Large Language Models (LLMs) by retrieving relevant external knowledge, thereby strengthening the performance of Generative AI (GenAI) applications. While GenAI has seen use in software engineering [19], RAG extends its value to broader domains by combining parametric and non-parametric memory, effectively addressing the limitations of static knowledge bases [27]. Recent advances in retrieval-augmented LLMs enable real-time information retrieval, reducing hallucinations and improving response reliability [7].
The foundational work by Lewis et al. [14] established RAG as a standard for tasks like question answering and knowledge retrieval. Early frameworks such as REALM and RAG demonstrated the benefits of combining dense retrieval with language generation for open-domain tasks [10]. However, most later research has focused on improving retrieval architectures and reducing hallucinations, often evaluated only on clean, English-centric benchmarks [6]. There remains limited exploration of domain-specific, multilingual, or real-world deployments, which this paper addresses through the development and evaluation of RAG systems focused on retrieval quality and system design.
@@ 1 Introduction
This study investigates how RAG systems can be engineered and evaluated in real-world contexts, focusing on system design, domain-specific applications, user evaluation, and lessons for future practice. Guided by these aims, this paper addresses the following Research Questions (RQs):
  • RQ1: How can RAG systems be designed and developed to address real-world system needs across diverse application domains, and what are the lessons learned from engineering RAG systems for real-world applications?
@@
To address these research questions, we developed five domain-specific RAG systems and evaluated them through a web-based user study with 100 participants. The evaluation focused on usability, retrieval relevance, transparency, and other user-centred factors. Full methodological details are provided in Section 3.
The contributions of this paper are as follows: end-to-end development and deployment of RAG systems for multilingual, domain-specific applications; user-centred evaluation demonstrating real-world performance across usability and accuracy metrics; practical engineering insights to guide the design of reliable and maintainable RAG pipelines; and system-level considerations for integrating RAG into real-world AI-based software, contributing to software engineering practice.
Paper Structure: Section 2 reviews related work. Section 3 outlines the study design. Section 4 details the system implementation. Section 5 presents the evaluation. Section 6 highlights key lessons learned. Section 7 discusses limitations, and Section 8 offers conclusions and future directions.
2 Related Work
RAG improves the factual accuracy and contextual relevance of LLMs by incorporating real-time external information, making it especially valuable for complex tasks such as question answering, legal reasoning, and summarization [20]. Recent work has demonstrated RAG’s utility in taxonomy-driven dataset design [16], token-efficient document handling [20], and multimodal applications that combine text and images via Vision-Language Models (VLMs) such as VISRAG [29]. Despite these advances, OCR noise remains a limiting factor in retrieval fidelity [32]. Ongoing research addresses this by refining dataset construction [16], tackling architectural scalability [5], improving query-document alignment through prompt engineering [34], and applying speculative retrieval to boost performance in multimodal settings [33].
RAG has been applied in software engineering for code understanding and developer tasks, e.g., StackRAG [1] enhances developer assistance using Stack Overflow, and CodeQA [2] applies LLM agents with retrieval augmentation. Ask-EDA [22] reduces hallucinations in Electronic Design Automation via hybrid retrieval. In industry, Khan et al. [12] address PDF-focused retrieval, while Xiaohua et al. [26] propose re-ranking and repacking for pipeline optimization. In healthcare, MEDGPT [24] extracts structured diagnostic insights, Path-RAG [18] improves pathology image retrieval, Alam et al. [3] introduce multi-agent retrieval for radiology reports, and Guo et al. [9] present LightRAG for graph-based precision retrieval.
@@ 2 Related Work
The rise of autonomous AI agents has further improved RAG by enabling self-directed reasoning, adaptive retrieval, and memory persistence. Wang et al. [25] survey LLM-driven agent architectures, while Liu et al. [17] benchmark multi-turn reasoning through Agent-bench. Agent-tuning by Zeng et al. [30] enhances instruction tuning for retrieval-based decisions. Singh et al. [23] categorize Agentic RAG into single-agent, multi-agent, and graph-based designs, highlighting dynamic tool use. On the retrieval side, Yan et al. [28] introduce CRAG to reduce hallucinations using confidence-based filtering, and Li et al. [15] improve precision through contrastive in-context learning and focus-mode filtering, strengthening RAG’s reliability in complex scenarios.
Conclusive Summary: While RAG continues to advance, challenges in retrieval accuracy, reliability, and scalability persist [13]. Despite progress in hybrid strategies [21], autonomous agents [23], and correction techniques [28, 15, 31], domain-specific evaluation remains limited. This paper evaluates five RAG systems and offers practical development and deployment insights.
3 Study Design
Figure 1 provides an overview of the methodological steps, detailing the phased implementation of five domain-specific RAG systems, a user-centred evaluation with 100 participants, and the synthesis of lessons learned across technical, operational, and ethical dimensions.
@@ 3 Study Design
3.1 Implementing RAG Systems
@@
3.1.2 System Design.
@@
3.2 System Evaluation Method
@@
After using the systems, participants completed a standardized survey covering six criteria: Ease of Use, Relevance of Info, Transparency, System Responsiveness, Accuracy of Answers, and Recommendation. The survey included both Likert-scale questions and open-ended feedback. This approach provided both quantitative ratings and qualitative insights into system performance. We reviewed the open-ended feedback to identify common themes in participants’ experiences. We also referred to development notes taken throughout the project. These helped us recognize recurring issues and informed the lessons described in Section 6.
4 Systems Implementation
This section outlines the implementation of five RAG-based systems for real-world deployment across diverse domains. Figure 2 presents the system architecture, showing user interaction, domain-specific data processing, vector storage, retrieval, and response generation with LLMs.
@@ 4 Systems Implementation
    Kankaanpää City AI: This system enhances transparency of government records. It processes over 1,000 PDFs from 2023–2024, indexing them in FAISS for accurate retrieval of policy documents. The system uses the embedding model text-embedding-ada-002 to convert documents into vector representations, and gpt-4o-mini as the LLM to generate context-aware responses. This setup allows users to search and access municipal decisions, infrastructure projects, and public policies with ease.
@@
    Disarm RAG: It is designed to deliver real-time insights into cyber threats, and forensic investigations. The system uses LLaMA 2-uncensored via Ollama to enable open access to cybersecurity knowledge, and is hosted on a secure server at CSC [^6] https://research.csc.fi/cloud-computing/ (Finnish IT Center for Science) to ensure full data privacy. The system integrates red team techniques (e.g., phishing, deep-fake disinformation, privilege escalation) and blue team strategies (e.g., bot detection, misinformation control, network forensics), grounded in the Disarm Framework. It supports queries such as “How would you create a deep-fake to discredit a public figure?” and “What are the latest techniques for bypassing multi-factor authentication (MFA)?”, as well as defensive questions like “How would you detect a disinformation campaign early on?” and “What are effective countermeasures against deep-fake-based phishing attacks?”.
Figure 2: System architecture of five RAG-based systems, showing data processing, vector storage, retrieval, and LLM-based response generation.
@@
All developed systems comply with GDPR and display source references, except Disarm RAG, where citations are hidden for security reasons.
5 Systems Evaluation
Understanding the real-world effectiveness of RAG-based systems requires moving beyond technical benchmarks to incorporate user-centred evaluation. We conducted a structured user study across five domain-specific deployments, capturing both system performance metrics and user perceptions of trust, relevance, and usability. This practical feedback offers a grounded view of system behaviour in real settings and highlights opportunities for targeted improvements.
5.1 Participant Demographics and RAG Orientation
@@
    AI-Generated vs. Manual Document Search: Participants exhibited a task-sensitive perspective on AI assistance. While a substantial majority (83%) preferred AI-generated responses depending on the nature of the task, only (9%) expressed a consistent preference for AI over manual methods. Conversely, (8%) favoured manual search regardless of context. These findings suggest that trust in RAG systems is not absolute but contingent—underscoring the importance of response relevance, transparency, and alignment with user intent.
    Figure 3: Participant profiles and interaction with the RAG systems.
@@
5.2 Survey Instrument and Case-wise Findings
@@
6 Lessons Learned
While developing and evaluating the five RAG systems, we encountered technical, operational, and ethical challenges. These lessons reflect practical insights drawn from our hands-on engineering work, system deployment in real-world domains, and user-centred feedback. Together, they offer guidance for building reliable, domain-adapted RAG systems that balance performance, compliance, and trustworthiness.
6.1 Technical Development
@@
  • FAISS Scalability Hits Limits: With large corpora (>>10k embeddings), FAISS latency increased noticeably. Metadata filtering by document type reduced search time.
@@ 6.2 Operational Factors
  • SQLite for Tracking User Interaction: We used SQLite to log user questions, responses, and ratings (e.g., in AgriHubi). This lightweight store helped identify system failures and understand user behaviour.
@@ 6.3 Ethical Considerations
7 Study Limitations
This study presents findings grounded in the design, deployment, and evaluation of five domain-specific RAG systems, but several limitations must be acknowledged. First, while our evaluation involved 100 participants across diverse roles including researchers, practitioners, and domain experts, approximately 20% of the sample consisted of students. Although these students had relevant technical or domain experience, their feedback may reflect differing expectations or usage behaviour compared to full-time professionals. This demographic distribution, while broad, could influence the generalizability of findings to strictly industrial settings. Additionally, participants rated the accuracy of system-generated answers, yet only 17% identified as domain experts. Thus, accuracy ratings provided by non-experts might not reliably reflect the factual correctness of the systems’ outputs.
@@ 8 Conclusion
Acknowledgements
This paper is based on research supported in part by the Synthetica project funded by the Research Council of Finland, the GENT project funded by the European Regional Development Fund, and the AgriHubi initiative supported by Pohjois-Pohjanmaan ELY-keskus (Centre for Economic Development, Transport and the Environment). The authors gratefully acknowledge this support. The authors declare no conflicts of interest related to this work.